The concentration of a hypothetical molecule, A5B6C2, is 943 mg/dL.   The molar masses of the hypothetical elements are in the table below. What is the molarity of this molecule in the solution?

Element Molar Mass (grams/mole)
A 10.852
B 1.0797
C 15.35
Answer #1

Molecular mass of A5B6C2 = (10.852 x 5) + (1.0797 x 6) + (15.35 x 2) = 54.26 + 6.4782 + 30.7 = 91.4382 g

It is given in the question that concentration is 943 mg/dl which means that 943 mg of solute in 1 dl of water.

No. of moles of solute = Given Mass / Molecular Mass = 943 mg / 91.4382 g

1 mg = 0.001 g

No. of moles of solute = Given Mass / Molecular Mass = 943 X 0.001 g / 91.4382 g = 0.0103 moles

Given volume = 1 dl = 0.1 litre

Molarity = Number of moles of solute / Volume of solution in litre = 0.0103 / 0.1 mol/l = 0.103 mol/l